OpenStudy (ammarah):

Find all locations (x,y) on the graph of f(x)= x^2 +4x+3 where slope =10

OpenStudy (ammarah):

@girondasearch @girondasearch @zepdrix

zepdrix (zepdrix):

Using calculus techniques?

OpenStudy (ammarah):

yes!

OpenStudy (ammarah):

I'm in precalc

zepdrix (zepdrix):

The derivative gives us slope at a particular point, so we're looking for x values such that\[\large\rm f'(x)=10\]Do you understand how to find f'(x)?

OpenStudy (photon336):

calculate the derivative and set it equal to 10 :)

OpenStudy (ammarah):

the slope is given which is 10

OpenStudy (photon336):

so they're asking the points where the slope i.e. derivative is 10

OpenStudy (photon336):

so @Ammarah do what @zepdrix said, but first find the derivative for us

OpenStudy (ammarah):

do I use f(x +h)-f(x)/h?

OpenStudy (photon336):

there is a much easier way

OpenStudy (ammarah):

can someone help me solve and graph this please

OpenStudy (ammarah):

do i plug in points?

OpenStudy (photon336):

to take the derivative of a polynomial here's your formula you need to use \[\frac{ d }{ dx }x^{n} = nx^{n-1}\]

OpenStudy (ammarah):

what is n

OpenStudy (photon336):

so n is the exponent

OpenStudy (photon336):

\[x^{2}\] say if I had that can you tell me what n would be?

OpenStudy (ammarah):

2

OpenStudy (photon336):

so using the formula above what would be the derivative of \[x^{2}\]

OpenStudy (ammarah):

idk

OpenStudy (ammarah):

is ther another way? I never learned this formula

OpenStudy (photon336):

Just try, \[\frac{ d }{ dx }x^{n} = nx^{n-1}\]

OpenStudy (ammarah):

like can i use point slope form or plug in points for x?

OpenStudy (photon336):

you have to follow me, because this is the fastest way

OpenStudy (photon336):

we take the exponent and we subtract 1 right? so in x^{2} the exponent is 2. 2-1 = 1 then 2 goes in front of the x \[\frac{ d }{ dx } x^{2} = 2x \]

OpenStudy (photon336):

@Ammarah can you see what the pattern is?

OpenStudy (ammarah):

yes

OpenStudy (ammarah):

then what do I do?

zepdrix (zepdrix):

Oh you're using the limit definition of derivative?

zepdrix (zepdrix):

\[\large\rm \lim_{h\to0}\frac{f(x+h)-f(x)}{h}\]yes?

OpenStudy (ammarah):

so I would just plug that in? how do I find all the locations?

zepdrix (zepdrix):

Yes, that will give you the function which determines the locations. Make sure you understand how to plug in the pieces, it can be a little tricky. \[\large\rm f(\color{orangered}{x})=(\color{orangered}{x})^2+4(\color{orangered}{x})+3\]So the other piece in the numerator is going to be,\[\large\rm f(\color{orangered}{x+h})=(\color{orangered}{x+h})^2+4(\color{orangered}{x+h})+3\]

zepdrix (zepdrix):

\[\large\rm \frac{f(x+h)-f(x)}{h}\]So we plug in our numerator,\[\large\rm \frac{(x+h)^2+4(x+h)+3-(x^2+4x+3)}{h}\]Distribute all the negatives and stuff, expand out the square,\[\large\rm \frac{x^2+2xh+h^2+4x+4h+3-x^2-4x-3}{h}\]Combine like-terms,\[\large\rm \frac{2xh+h^2+4h}{h}\]Divide out an h,\[\large\rm 2x+h+4\]Then the calculus part comes in, this idea of "limits". I dunno if you've been introduced to this? You care about what happens to this function as h (the space between your two points in your slope formula) goes to zero. What that means for us here, is that we can directly plug 0 in for h.

zepdrix (zepdrix):

\[\large\rm f'(x)=2x+4\]That gives us our derivative function, this gives us the slope of the original function at any point. We care about when the slope is 10,\[\large\rm 10=2x+4\]

zepdrix (zepdrix):

Sorry if I took some of the fun away, doing the whole problem like that, you seemed a bit confused though :d

OpenStudy (ammarah):

thank you so much! so this equation basically defines all the locations on the graph, like if I plugged in a value for x and y?

zepdrix (zepdrix):

Here is what our function looks like.

zepdrix (zepdrix):

the derivative function tells us how the function is changing, how steeply the function is growing at a particular moment in time. So for example, here you can see that at x=-2, the function is not growing at all, it's stationary, while at x=1, it's very very steep, maybe a slope value of 5 or 6.